OpenStudy (anonymous):

Need help approaching this question, just started taking physics this year. A train accelerates steadily from 4 ms^-1 to 20ms^-1 in 100s. From its initial and final velocities, calculate the average velocity of the train.

OpenStudy (abhisar):

Hello @tombraider5000 ! \(\sf \huge Average~Velocity=\frac{Total~Displacement}{Time~Taken}\)

OpenStudy (abhisar):

STEP 1 Can you calculate the acceleration by using the eqaution v=u+at ?

OpenStudy (abhisar):

STEP 2 Now calculate the total displacement by using the equation s=ut+\(\sf\frac{1}{2}a\times t^2\)

OpenStudy (abhisar):

STEP 3 Finally calculate the average velocity